OpenStudy (jack1):

need help expanding out the brackets please: (4x^3 +4xy) times (1 - 2x^2 - 2y)^-1

OpenStudy (jim766):

if the 2nd polynomial is all to the -1 power it would look like this \[(4x^3+4xy)*\frac{ 1 }{ (1-2x^2-2y) }\]

OpenStudy (jack1):

yep, but how do i expand it out?

OpenStudy (jim766):

\[\frac{ 4x^3+4xy }{ 1-2x^2-2y }\]

OpenStudy (jack1):

Im doing ODE's, it's currently: y' = (4x^3 +4xy) /(1 - 2x^2 - 2y) need it in the form: y' + (P(x)) y = Q(x)

OpenStudy (jack1):

@Peter14 lil help if you got a sec?

OpenStudy (anonymous):

see...

OpenStudy (anonymous):

multiply the whole thing by (1−2x^2−2y)

OpenStudy (tkhunny):

y' = (4x^3 +4xy) /(1 - 2x^2 - 2y) You may wish to observe this: \(y' = \dfrac{2x\cdot(2x^{2}+2y)}{1 - (2x^{2} + 2y)}\)

OpenStudy (jack1):

so y ' = 2x (2x^2 +2y)/1 - (2x (2x^2 +2y)/(2x^2 +2y)?

OpenStudy (jack1):

which equals : y ' = 2x (2x^2 +2y) - 2x ...?

OpenStudy (tkhunny):

Have you considered substitution from your studies of integration? Take a look at the derivative (dy/dx) of \(ln(1-(2x^{2} + 2y))\). I'm not saying this leads directly to a solution. Just think on it.

OpenStudy (jack1):

no, sorry at the moment i'm only comfortable with solving from : y' + (P(x)) y = Q(x) and finding the integrating factor

OpenStudy (jack1):

thanks for the help tho